help me set up equation and solve.
A farmer grows only sunflowers and flax on his 240-acre farm. This year, he planted 80 more acres of sunflower than of flax. How many acres of each crop has he planted?

Answer by LinnW(1048)

set f = acres of flax , s = acres of sunflowers
f + s = 240
s = f + 80
substitute (f + 80) for s in f + s = 240
f + (f + 80) = 240
2f + 80 = 240
add -80 to each side
2f = 160
divide each side by 2
f = 80
since s = f + 80 ,
s = 80 + 80
s = 160